OpenStudy (anonymous):

differential equations e^xdx-(1+e^-x)^2dy=0

OpenStudy (anonymous):

It looks separable!

OpenStudy (anonymous):

can you help me with it? please

OpenStudy (anonymous):

Switch the term with x, dx to the right side!

OpenStudy (anonymous):

Can you do it?

OpenStudy (anonymous):

Just isolate dy Then integrate both side to get y !

OpenStudy (anonymous):

i am trying

OpenStudy (anonymous):

i dont get how i should integrate the right side

OpenStudy (anonymous):

e^x dx-(1+e^-x)^2dy=0 (1+e^-x)^2dy = e^xdx => dy = e^xdx / (1+e^-x)^2

OpenStudy (anonymous):

Agree?

OpenStudy (anonymous):

yes, and what then? do we do it by partial? :/ i am so lost i followed u until there

OpenStudy (anonymous):

Since you want y! -> Int dy = int e^xdx / (1+e^-x)^2 Agree?

OpenStudy (anonymous):

oh we take ln on both sides?

OpenStudy (anonymous):

Integrate dy => y ( not ln)

OpenStudy (anonymous):

yes this is where i am getting stuck at i can seem to find away to integrate e^x/(1-E^-x)^2 dx

OpenStudy (anonymous):

i cant*

OpenStudy (anonymous):

a way*

OpenStudy (anonymous):

should i open the square?

OpenStudy (anonymous):

Let u = 1 + e^-x -> du = ?

OpenStudy (anonymous):

e^-x?

OpenStudy (anonymous):

-e^-x*

OpenStudy (anonymous):

du = - e^-x dx

OpenStudy (anonymous):

Is it okie for you to continue?

OpenStudy (anonymous):

but in numerator its e^x not e^-x ? how am i suppose to substiute du?

OpenStudy (anonymous):

=> -du = e^-x dx

OpenStudy (anonymous):

Oh, let me think!

OpenStudy (anonymous):

yes but still on top it is e^x and my (-)du is e^-x? isnt du suppose to by e^+x dx?

OpenStudy (anonymous):

in order to substitute

OpenStudy (anonymous):

I'm thinking of - 1/ du = e^x dx

OpenStudy (anonymous):

um idk if its gunna help :/

OpenStudy (anonymous):

Nope, so in the denominator: (1 + 1/e^x )^2 Let u = e^x

OpenStudy (anonymous):

oooooooooooohhhh lemme try that.. thankyou sooo muchhhhh..

OpenStudy (anonymous):

My pleasure :D

OpenStudy (anonymous):

i end up with integral of 1/(1+1/u)^2 du?

OpenStudy (anonymous):

= Int ( u^2 du / ( u +1 ) ^2

OpenStudy (anonymous):

huh? o.O

OpenStudy (anonymous):

1 + 1/u = ( u +1 )/ u

OpenStudy (anonymous):

umm hold ... you substituted the top with "u" but where did "du" go? du is e^x dx

OpenStudy (anonymous):

From (1 + 1/ e^x )^2, take u = e^x => du = e^xdx ( on the top)

OpenStudy (anonymous):

ohh okay so its integral u^2/(u+1)^2 du

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

now we make another substitution?

OpenStudy (anonymous):

no non o

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

nvm???

OpenStudy (anonymous):

nevermind lol

OpenStudy (anonymous):

I thought some technical term :P

OpenStudy (anonymous):

lol haha, hey i made another substitution, is that right?

OpenStudy (anonymous):

Seem like it !

OpenStudy (anonymous):

i took v=1-u: dv=du and u=v-1

OpenStudy (anonymous):

and then i opened the square on the top

OpenStudy (anonymous):

Should we use partial ?

OpenStudy (anonymous):

(v-1)^2/v^2= integral v^2-2v+1/v^2; v^2/v^2-2v^v+1/v^2 dv?

OpenStudy (anonymous):

last part v^2/v^2-2v/v^2+1/v^2 dv

OpenStudy (anonymous):

Let see if my answer is same as yours!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

e^x+3-1/2(e^x+1)^2+c idk :/

OpenStudy (anonymous):

I have 3 parts, but stuck at the second: Int ( udu / (u + 1)^2 )

OpenStudy (anonymous):

omg i hate calculus, its fine chloriphyllthankyou very much i got some idea. you helped me ALOT.. thankyou again.. GOD bless you.. i got some more problems. to do also.. :/

OpenStudy (anonymous):

Okie, I'll get back to it later! Just move on to the next one :(

OpenStudy (anonymous):

Now I hate e^-x awfully !

OpenStudy (anonymous):

lol, m sorry :/

OpenStudy (anonymous):

Just move to the next one! I'm pretty sure some higher level OS person will come to rescue!

OpenStudy (anonymous):

The last choice is input into wolfram, though!

OpenStudy (anonymous):

btw do you how to do those surface area problems? the one with this formula= integral 2pi f(x) sqrt(1+(f ' (x))^2 ?

OpenStudy (anonymous):

Just open the new post!

OpenStudy (anonymous):

okie dokie

OpenStudy (anonymous):

@bmp Can you take a look at this post, see if you can help?

OpenStudy (anonymous):

Do you need help with what integral? Or with the DE itself?

OpenStudy (anonymous):

The integral

OpenStudy (anonymous):

udu / (u + 1)^2 This?

OpenStudy (anonymous):

pleaseee

OpenStudy (anonymous):

dy = e^xdx / (1+e^-x)^2

OpenStudy (anonymous):

I'm stuck with this integral !

OpenStudy (anonymous):

Make u = e^x, du = e^x dx, then the integral becomes: \[ \int \frac{du}{(1/u + 1)^2} \]From there, it's a long division, and integrating.

OpenStudy (anonymous):

Should we convert into Int ( u^2 du / (u +1)^2 )

OpenStudy (anonymous):

Sorry, my browser crashed twice. Not exactly that, notice that we get e^(-x) on the bottom, and we have e^x on top. Not the same thing.

OpenStudy (anonymous):

It should ended up with what I posted above, because if u = e^x, then 1/u = e^-x

OpenStudy (anonymous):

I did get to that result, but I don't know how to proceed further!

OpenStudy (anonymous):

Do a long division of that, you should get something like 1 - 2/(u + 1) + 1/(u+1)^2. :-)

OpenStudy (anonymous):

Expand the denominator out and do partial fractions. I think that should do it too.

OpenStudy (anonymous):

I seem go around with the result :(

OpenStudy (anonymous):

Maybe I should have made myself clearer, 1/(1/x + 1)^2 = x^2/(x^2 + 1)^2 I think it's easier to see it now (the division, I mean).

OpenStudy (anonymous):

Yep, that's what I came up! Int ( u^2 du / ( u +1 ) ^2 )

OpenStudy (anonymous):

Do the polynomial long division now. What is x^2/(x^2 + 1)^2? :-)

OpenStudy (anonymous):

Or partial fractions, both work just fine :-)

OpenStudy (anonymous):

I thought the inside is ( x +1 )^2

OpenStudy (anonymous):

Ah, my bad. You are correct, it's only (x+1)^2.

OpenStudy (anonymous):

A ( x +1) + ( Bx + C ) = x^2

OpenStudy (anonymous):

Yup, albeit I am a bit rusty with my partial fractions :-(

OpenStudy (anonymous):

I think x = -1 seems a good candidate.

OpenStudy (anonymous):

Does it look right? A + B = 0 A + B + C = 0

OpenStudy (anonymous):

Hmmm, yeah, I got a similar problem here.

OpenStudy (anonymous):

Oh, I set up the partial wrong!

OpenStudy (anonymous):

x^2 / (x + 1)^2 = A/ ( x+1) + B ( x + 1) ^2

OpenStudy (anonymous):

-> A ( x+1) + B = x^2

OpenStudy (anonymous):

Shouldnt a +C show up also? The long division yielded a +1 elsewhere.

OpenStudy (anonymous):

A C/(x+1)^2, that is.